OpenStudy (anonymous):

find a ratio equivalent to 7/8

OpenStudy (anonymous):

theres lot of infinite number of a ratio equivalent to 7/8 example 7/8 *2/2 = 14/16 , 7/8 * 3/3 = 21/24 , 7/8 *4/4=28/32 etc...etc....

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

YW good luck now and have fun :D